# Test-Time Compute: from System-1 Thinking to System-2 Thinking

### v2

@@ 3.2 Modifying the Input
First, empirical studies (Liu et al., 2022) show that the more similar the demonstrations are to the test sample, the better the ICL performance. Therefore, retrieval models like BM25 and SentenceBERT are used to retrieve demonstrations semantically closest to the test sample and rank them in descending order of similarity (Qin et al., 2024a; Luo et al., 2023a). To improve the accuracy of demonstration retrieval, Rubin et al. (2022) and Li et al. (2023b) specifically train the demonstration retriever by contrastive learning. Then, as researchers delve deeper into the mechanisms of ICL, ICL is considered to conduct implicit gradient descent on the demonstrations (Dai et al., 2023). Therefore, from the perspective of training data, demonstrations also need to be informative and diverse (Su et al., 2022; Li and Qiu, 2023). Wang et al. (2023c) view language models as topic models and formulate the demonstration selection problem as solving a Bayesian optimal classifier. Additionally, the ordering of examples is another important area for improvement. Lu et al. (2022) and Wu et al. (2023) use information theory as a guide to select the examples with maximum local entropy and minimum description length for ranking, respectively. Scarlatos and Lan (2024) and Zhang et al. (2022b) consider the sequential dependency among demonstrations, and model it as a sequential decision problem and optimize demonstration selection and ordering through reinforcement learning.
@@
Another line of work (Chen et al., 2023; Lyu et al., 2023; Kim et al., 2022; Zhang et al., 2023d) argues that in practice, combining a limited set of externally provided examples may not always be the optimal choice. LLMs can leverage their generative and annotation capabilities to create better demonstrations. DAIL (Su et al., 2024) constructs a demonstration memory, storing previous test samples and their predictions as candidate demonstrations for subsequent samples. DAWN-ICL (Tang et al., 2024a) further models the traversal order of test samples as a planning task and optimizes it by the Monte Carlo tree search (MCTS).
3.3 Editing the Representation
@@
For generative LLMs, some works have found that the performance bottleneck is not in encoding world knowledge, but in the large gap between the information in intermediate layers and the output. During the inference phase, editing the representation can help externalize the intermediate knowledge into the output. PPLM (Dathathri et al., 2020) performs gradient-based representation editing under the guidance of a small language model to control the style of outputs. ActAdd (Turner et al., 2024) selects two semantically contrastive prompts and calculates the difference between their representations as a steering vector, which is then added to the residual stream. Representation editing based on contrastive prompts has demonstrated its effectiveness in broader scenarios, including instruction following (Stolfo et al., 2024), alleviating hallucinations (Li et al., 2023a; Arditi et al., 2024), reducing toxicity (Liu et al., 2024b; Lu and Rimsky, 2024) and personality (Cao et al., 2024). SEA (Qiu et al., 2024b) projects representations onto directions with maximum covariance with positive prompts and minimum covariance with negative prompts. They also introduce nonlinear feature transformations, allowing representation editing to go beyond linearly separable representations. Scalena et al. (2024) conduct an in-depth study on the selection of steering intensity. They find that applying a gradually decreasing steering intensity to each output token can improve control over the generation without compromising quality.
Figure 3: Illustration of feedback modeling, search strategies and improvement training in test-time reasoning.
3.4 Calibrating the Output
@@
Using external information to calibrate the model’s output distribution is also an efficient yet effective test-time adaptation method (Khandelwal et al., 2020). AdaNPC (Zhang et al., 2023c) designs a memory pool to store training data. During inference, given a test sample, AdaNPC recalls $k$ samples from the memory pool and uses a $k$NN classifier to predict the test sample. It then stores the test sample and its predicted label in the memory pool. Over time, the sample distribution in the memory pool gradually aligns with the test distribution. In NLP, the most representative application of such methods is $k$NN machine translation ($k$NN-MT). $k$NN-MT (Khandelwal et al., 2021) constructs a datastore to store contextual representations and their corresponding target tokens. During translation inference, it retrieves the $k$-nearest candidate tokens from the datastore based on the decoded context and processes them into probabilities. Finally, it calibrates the translation model’s probability distribution by performing a weighted fusion of the model’s probabilities and the retrieved probabilities. $k$NN-MT has demonstrated superior transferability and generalization compared to traditional models in cross-domain and multilingual MT tasks. Subsequent studies have focused on improving its performance and efficiency (Wang et al., 2022a; Zhu et al., 2023b; You et al., 2024) or applying its methods to other NLP tasks (Wang et al., 2022b; Bhardwaj et al., 2023).
@@
4 Test-time Reasoning for System-2 Thinking
@@ 4.1 Feedback Modeling
Score-based Discriminative Cobbe et al. (2021) Math classification ORM; human annotated data
@@
Acemath (2025b) Math list-wise Bradley-Terry ORM; sampling training data from multiple LLMs
Lightman et al. (2024) Math classification PRM; human annotated data
Math-shepherd (2024i) Math classification PRM; annotating processes via MCTS
@@
Zhang et al. (2025) Math classification/regression PRM; annotating processes via MCTS and LLM-as-a-judge
@@
Implicit PRM (2024b) Math implicit reward modeling PRM; training PRMs with outcome labels
Generative GenRM (2024g) Math SFT PRM; synthesizing critique data from external LLMs
Critic-RM (2024b) General SFT & Bradley-Terry ORM; synthesizing and filtering critique data via self-critique
CLoud (2024) General SFT & Bradley-Terry ORM; synthesizing data from external LLMs and self-critique
@@
Verbal-based Training-free LLM-as-a-Judge (2023) General - Designing system instructions to mitigate biases
@@
BSM (2024) General - Dividing into multiple criteria and then merging
@@
Training-dependent Shepherd (2023b) General SFT Collecting data from human annotation and the Internet
Auto-J (2024b) General SFT Collecting data from GPT-4
@@
Prometheus (2024c) General SFT Training single and pairwise models and then merging them
@@
EvalPlanner (2025) General DPO Planing evaluation processes and then evaluating
Table 1: Overview of feedback modeling methods.
Score-based feedback, also known as the verifier, aims to score generated results, evaluating their alignment with ground truth or human cognitive processes. Its training process is typically similar to the reward model in RLHF, using various forms of feedback signals and modeling it as a classification (Cobbe et al., 2021) or rank task (Bradley and Terry, 1952; Yuan et al., 2024a; Hosseini et al., 2024). In reasoning tasks, verifiers are mainly divided into two categories: outcome-based (ORMs) and process-based verifiers (PRMs). ORMs (Cobbe et al., 2021) use the correctness of the final CoT result as training signals. Liu et al. (2025b) provide a detailed recipe for training a strong ORM. In contrast, PRMs (Uesato et al., 2022; Lightman et al., 2024; Zhang et al., 2024e) are trained based on the correctness of each reasoning step. Compared to ORMs, PRMs not only can evaluate intermediate reasoning steps but also assess the entire reasoning process more precisely. However, PRMs require more human effort to annotate feedback for the intermediate steps. Math-Shepherd (Wang et al., 2024i) and OmegaPRM (Luo et al., 2024) utilize MCTS algorithm to collect high-quality process supervision data automatically. Zhang et al. (2025) utilize critic models to evaluate process annotations collected by MCTS, filtering out low-quality data to improve the training effectiveness of PRMs. Setlur et al. (2024) argue that PRMs should evaluate the advantage of each step for subsequent reasoning rather than focusing solely on its correctness. They propose process advantage verifiers (PAVs) and efficiently construct training data through MCTS. Furthermore, Lu et al. (2024) and Yuan et al. (2024b) notice that ORMs implicitly model the advantage of each step, leading them to automatically annotate process supervision data using ORMs or directly train PRMs on outcome labels, respectively.
Score-based feedback modeling overlooks the generative capabilities of LLMs, making it difficult to detect fine-grained errors. Thus, recent works propose generative score-based verifiers (Ankner et al., 2024; Ye et al., 2024). GenRM (Zhang et al., 2024g) leverages instruction tuning to enable the verifier to answer ‘Is the answer correct (Yes/No)?’ and uses the probability of generated ‘Yes’ token as the score. GenRM can also incorporate CoT, allowing the verifier to generate the corresponding rationale before answering ‘Yes’ or ‘No’. Critic-RM (Yu et al., 2024b) jointly trains the critique model and the verifier. During inference, the verifier scores according to answers and verbal-based feedback generated by the critique model.
Verbal-based Feedback
Although the verifier can accurately evaluate the correctness of generated answers or steps, it lacks interpretability, making it unable to locate the specific cause of errors or provide correction suggestions. Verbal-based feedback, also referred to critic, fully leverages the LLM’s instruction-following ability. By designing specific instructions, it can perform pairwise comparisons, evaluate answers from multiple dimensions, and even provide suggestions for revision in natural language. Powerful closed-source LLMs, such as GPT-4 and Claude, are effective critics. They can perform detailed and controlled assessments of generated texts, such as factuality, logical errors, coherence, and alignment, with high consistency with human evaluations (Wang et al., 2023a; Luo et al., 2023b; Liu et al., 2023; Chiang and Lee, 2023). However, they still face biases such as length, position, and perplexity (Bavaresco et al., 2024; Wang et al., 2024h; Stureborg et al., 2024). LLM-as-a-Judge (Zheng et al., 2023) carefully designs system instructions to mitigate the interference of biases. BSM (Saha et al., 2024) evaluates based on multiple criteria and then merges them.
To obtain cheaper verbal-based feedback, open-source LLMs can also serve as competitive alternatives through supervised fine-tuning  (SFT) (Wang et al., 2024m; Zhu et al., 2023a; Liang et al., 2024c; Paul et al., 2024). Shepherd (Wang et al., 2023b) collects high-quality training data from human annotation and online communities to fine-tune an evaluation model. Auto-J (Li et al., 2024b) collects queries and responses from various scenarios and designs evaluation criteria for each scenario. GPT-4 then generates critiques of the responses based on these criteria and distills its critique ability to open-source LLMs. Prometheus (Kim et al., 2024b, c) designs more fine-grained evaluation dimensions. It trains a single evaluation model and a pairwise ranking model separately, then unifies them into one LLM by weight merging. To reduce reliance on human annotations and external LLMs, Wang et al. (2024j) propose a self-training method: the critique model generates positive and negative responses, then collects critique data via rejection sampling to perform iterative finetuning. Building on self-training, EvalPlanner enables (Saha et al., 2025) the critique model to plan evaluation processes and criteria, conduct critiques based on these, and then collect positive and negative samples to improve the critique model via DPO (Rafailov et al., 2023).
@@
Repeat Sampling Majority voting CoT-SC (2023d) Math, QA self-consistency $\checkmark$
PROVE (2024) Math compiler $\checkmark$
@@
DiVeRSe (2023c) Math PRM
Knockout (2025a) Math critic $\checkmark$
Self-correction Human feedback NL-EDIT (2021) Semantic parsing Human
FBNET (2022) Code Human
External tools DrRepair (2020) Code compiler
Self-debug (2024d) Code compiler $\checkmark$
CRITIC (2024) Math, QA, Detoxifying text-to-text APIs $\checkmark$
External models REFINER (2024) Math, Reason critic model
Shepherd (2023b) QA critic model
Multiagent Debate (2023b) Math, Reason multi-agent debate $\checkmark$
MAD (2024b) Translation, Math multi-agent debate $\checkmark$
Intrinsic feedback Self-Refine (2023) Math, Code, Controlled generation self-critique $\checkmark$
Reflexion (2023) QA self-critique $\checkmark$
RCI (2023) Code, QA self-critique $\checkmark$
Tree Search Uninformed search ToT (2023) Planing, Creative writing self-critique $\checkmark$
@@
Heuristic search RAP (2023) Planing, Math, Logical self-critique $\checkmark$
TS-LLM (2024b) Planing, Math, Logical ORM
rStar (2024b) Math, QA multi-agent consistency $\checkmark$
ReST-MCTS* (2024a) Math, QA PRM
Table 2: Overview of search strategies.
4.2 Search Strategies
4.2.1 Repeated Sampling

### v3

@@ 3.2. Modifying the Input
First, empirical studies (Liu et al., 2022) show that the more similar the demonstrations are to the test sample, the better the ICL performance. Therefore, retrieval models like BM25 and SentenceBERT are used to retrieve demonstrations semantically closest to the test sample and rank them in descending order of similarity (Qin et al., 2024b; Luo et al., 2023b). To improve the accuracy of demonstration retrieval, Rubin et al. (2022) and Li et al. (2023b) specifically train the demonstration retriever by contrastive learning. Then, as researchers delve deeper into the mechanisms of ICL, ICL is considered to conduct implicit gradient descent on the demonstrations (Dai et al., 2023). Therefore, from the perspective of training data, demonstrations also need to be informative and diverse (Su et al., 2022; Li and Qiu, 2023). Wang et al. (2023d) view language models as topic models and formulate the demonstration selection problem as solving a Bayesian optimal classifier. Additionally, the ordering of examples is another important area for improvement. Lu et al. (2022) and Wu et al. (2023) use information theory as a guide to select the examples with maximum local entropy and minimum description length for ranking, respectively. Scarlatos and Lan (2024) and Zhang et al. (2022a) consider the sequential dependency among demonstrations, and model it as a sequential decision problem and optimize demonstration selection and ordering through reinforcement learning.
@@
Another line of work (Chen et al., 2023; Lyu et al., 2023; Kim et al., 2022; Zhang et al., 2023d) argues that in practice, combining a limited set of externally provided examples may not always be the optimal choice. LLMs can leverage their generative and annotation capabilities to create better demonstrations. DAIL (Su et al., 2024) constructs a demonstration memory, storing previous test samples and their predictions as candidate demonstrations for subsequent samples. DAWN-ICL (Tang et al., 2024) further models the traversal order of test samples as a planning task and optimizes it by the Monte Carlo tree search (MCTS).
3.3. Editing the Representation
@@
For generative LLMs, some works have found that the performance bottleneck is not in encoding world knowledge, but in the large gap between the information in intermediate layers and the output. During the inference phase, editing the representation can help externalize the intermediate knowledge into the output. PPLM (Dathathri et al., 2020) performs gradient-based representation editing under the guidance of a small language model to control the style of outputs. ActAdd (Turner et al., 2024) selects two semantically contrastive prompts and calculates the difference between their representations as a steering vector, which is then added to the residual stream. Representation editing based on contrastive prompts has demonstrated its effectiveness in broader scenarios, including instruction following (Stolfo et al., 2024), alleviating hallucinations (Li et al., 2023c; Arditi et al., 2024), reducing toxicity (Liu et al., 2024e; Lu and Rimsky, 2024) and personality (Cao et al., 2024). SEA (Qiu et al., 2024b) projects representations onto directions with maximum covariance with positive prompts and minimum covariance with negative prompts. They also introduce nonlinear feature transformations, allowing representation editing to go beyond linearly separable representations. Scalena et al. (2024) conduct an in-depth study on the selection of steering intensity. They find that applying a gradually decreasing steering intensity to each output token can improve control over the generation without compromising quality.
3.4. Calibrating the Output
@@
Using external information to calibrate the model’s output distribution is also an efficient yet effective test-time adaptation method (Khandelwal et al., 2020). AdaNPC (Zhang et al., 2023c) designs a memory pool to store training data. During inference, given a test sample, AdaNPC recalls $k$ samples from the memory pool and uses a $k$NN classifier to predict the test sample. It then stores the test sample and its predicted label in the memory pool. Over time, the sample distribution in the memory pool gradually aligns with the test distribution. In NLP, the most representative application of such methods is $k$NN machine translation ($k$NN-MT). $k$NN-MT (Khandelwal et al., 2021) constructs a datastore to store contextual representations and their corresponding target tokens. During translation inference, it retrieves the $k$-nearest candidate tokens from the datastore based on the decoded context and processes them into probabilities. Finally, it calibrates the translation model’s probability distribution by performing a weighted fusion of the model’s probabilities and the retrieved probabilities. $k$NN-MT has demonstrated superior transferability and generalization compared to traditional models in cross-domain and multilingual MT tasks. Subsequent studies have focused on improving its performance and efficiency (Wang et al., 2022a; Zhu et al., 2023a; You et al., 2024) or applying its methods to other NLP tasks (Wang et al., 2022b; Bhardwaj et al., 2023).
@@
4. Test-time Reasoning for System-2 Thinking
Figure 4. Illustration of feedback modeling, search strategies and improvement training in test-time reasoning.
@@ 4.1. Feedback Modeling
Score-based feedback, also known as the verifier, aims to predict scalar scores to evaluate the alignment of generated results with ground truth or human cognitive processes. Its training process is typically similar to the reward model in RLHF, using various forms of feedback signals and modeling it as a classification (Cobbe et al., 2021) or rank task (Bradley and Terry, 1952; Yuan et al., 2024a; Hosseini et al., 2024). In reasoning tasks, verifiers are primarily categorized into two types: outcome-based verifiers (ORMs) and process-based verifiers (PRMs). ORMs (Cobbe et al., 2021) use the correctness of the final CoT result as training signals. Liu et al. (2025b) provide a detailed recipe for training a strong ORM. In contrast, PRMs (Uesato et al., 2022; Lightman et al., 2024; Zhang et al., 2024f) are trained based on the correctness of each reasoning step. Compared to ORMs, PRMs cannot only evaluate intermediate reasoning steps but also assess the entire reasoning process more precisely. However, PRMs require more human effort to annotate feedback for the intermediate steps. Math-Shepherd (Wang et al., 2024g), OmegaPRM (Luo et al., 2024) and EpicPRM (Sun et al., 2025) utilize MCTS algorithm to collect high-quality process supervision data automatically. Zhang et al. (2025c) utilize critic models to evaluate process annotations collected by MCTS, filtering out low-quality data to improve the training effectiveness of PRMs. Setlur et al. (2024) argue that PRMs should evaluate the advantage of each step for subsequent reasoning rather than focusing solely on its correctness. They propose process advantage verifiers (PAVs) and efficiently construct training data through MCTS. Furthermore, Lu et al. (2024) and Yuan et al. (2024b) notice that ORMs implicitly model the advantage of each step, leading them to automatically annotate process supervision data using ORMs or directly train PRMs on outcome labels, respectively. Liu et al. (2025c) adaptively segment reasoning steps based on tokens with higher uncertainty, which is used to train better PRMs.
Generative-based Feedback
Although the verifier can evaluate the correctness of generated answers or steps, it lacks interpretability, making it unable to locate the specific cause of errors or provide correction suggestions. Generative-based feedback, also referred to critic, fully leverages the LLM’s generative and instruction-following ability (Ankner et al., 2024; Ye et al., 2024; Chiang et al., 2025; She et al., 2025). By designing specific instructions, it can perform pointwise or pairwise evaluation from multiple dimensions, and even provide suggestions for revision in natural language. Powerful closed-source LLMs, such as GPT-4 and Claude, are effective critics. They can perform detailed and controlled assessments of generated texts, such as factuality, logical errors, coherence, and alignment, with high consistency with human evaluations (Wang et al., 2023a; Luo et al., 2023a; Liu et al., 2023; Chiang and Lee, 2023). However, they still face biases such as length, position, and perplexity (Bavaresco et al., 2024; Wang et al., 2024f; Stureborg et al., 2024). LLM-as-a-Judge (Zheng et al., 2023) carefully designs system instructions to mitigate the interference of biases. BSM (Saha et al., 2024) evaluates based on multiple criteria and then merges them. Peng et al. (2025a) employ multi-agents to jointly evaluate answers’ factuality and instruction-following.
To obtain cheaper verbal-based feedback, open-source LLMs can also serve as competitive alternatives through supervised fine-tuning  (SFT) (Wang et al., 2024n; Zhu et al., 2023b; Liang et al., 2024b; Paul et al., 2024). Shepherd (Wang et al., 2023c) collects high-quality training data from human annotation and online communities to fine-tune an evaluation model. Auto-J (Li et al., 2024f) collects queries and responses from various scenarios and designs evaluation criteria for each scenario. GPT-4 then generates critiques of the responses based on these criteria and distills its critique ability to open-source LLMs. Prometheus (Kim et al., 2024c, d) designs more fine-grained evaluation dimensions. It trains a single evaluation model and a pairwise ranking model separately, then unifies them into one LLM by weight merging. To reduce reliance on human annotations and external LLMs, Wang et al. (2024d) propose a self-training method: the critique model generates positive and negative responses, then collects critique data via rejection sampling to perform iterative finetuning. Building on self-training, EvalPlanner enables (Saha et al., 2025) the critique model to plan evaluation processes and criteria, conduct critiques based on these, and then collect positive and negative samples to improve the critique model via DPO (Rafailov et al., 2023). (Yu et al., 2025b) carefully synthesize and filter data, enabling the base model to achieve strong critique abilities with only 40K samples for SFT and DPO training. GenRM (Zhang et al., 2024c) leverages instruction tuning to enable the verifier to answer ‘Is the answer correct (Yes/No)?’ and uses the probability of generated ‘Yes’ token as the score. GenRM can also incorporate CoT, allowing the verifier to generate the corresponding rationale before answering ‘Yes’ or ‘No’. ThinkPRM (Khalifa et al., 2025) evaluates each reasoning step with long CoT and requires only 1% of the process supervision data compared to discriminative PRMs. Critic-RM (Yu et al., 2024a) jointly trains the critic and the verifier.
@@
Score-based ORM Cobbe et al. (2021) Math classification ORM; human annotated data
@@
Acemath (Liu et al., 2025b) Math list-wise Bradley-Terry ORM; sampling training data from multiple LLMs
PRM Lightman et al. (2024) Math classification PRM; human annotated data
Math-shepherd (Wang et al., 2024g) Math classification PRM; annotating processes via MC estimation
@@
Zhang et al. (2025c) Math classification/regression PRM; annotating processes via MCTS and LLM-as-a-judge
@@
Implicit PRM (Yuan et al., 2024b) Math implicit reward modeling PRM; training PRMs with outcome labels
ASPRM (Liu et al., 2025c) Math, Code classification PRM; adaptive segment step; annotating processes via MC estimation
@@
Generative-based Training-free LLM-as-a-Judge (Zheng et al., 2023) General - Designing system instructions to mitigate biases
@@
BSM (Saha et al., 2024) General - Dividing into multiple criteria and then merging
@@
Training-based Shepherd (Wang et al., 2023c) General SFT Collecting data from human annotation and the Internet
@@
Prometheus (Kim et al., 2024d) General SFT Training single and pairwise models and then merging them
@@
EvalPlanner (Saha et al., 2025) General DPO Planing evaluation processes and then evaluating
GenRM (Zhang et al., 2024c) Math SFT PRM; synthesizing critique data from external LLMs
R-PRM (She et al., 2025) Math SFT & DPO PRM; synthesizing critique data from external LLMs
Critic-RM (Yu et al., 2024a) General SFT & Bradley-Terry ORM; synthesizing and filtering critique data via self-critique
CLoud (Ankner et al., 2024) General SFT & Bradley-Terry ORM; synthesizing data from external LLMs and self-critique
Table 1. Overview of feedback modeling methods.
@@
Repeat Sampling Majority voting CoT-SC (Wang et al., 2023b) Math, QA self-consistency $\checkmark$
PROVE (Toh et al., 2024) Math compiler $\checkmark$
@@
DiVeRSe (Li et al., 2023a) Math PRM
Knockout (Liu et al., 2025f) Math critic $\checkmark$
Self-correction Human feedback NL-EDIT (Elgohary et al., 2021) Semantic parsing Human
FBNET (Tandon et al., 2022) Code Human
External tools DrRepair (Yasunaga and Liang, 2020) Code compiler
Self-debug (Chen et al., 2024c) Code compiler $\checkmark$
CRITIC (Gou et al., 2024) Math, QA, Detoxifying text-to-text APIs $\checkmark$
External models REFINER (Paul et al., 2024) Math, Reason critic model
Shepherd (Wang et al., 2023c) QA critic model
Multiagent Debate (Du et al., 2023b) Math, Reason multi-agent debate $\checkmark$
MAD (Liang et al., 2024a) Translation, Math multi-agent debate $\checkmark$
Intrinsic feedback Self-Refine (Madaan et al., 2023) Math, Code, Controlled generation self-critique $\checkmark$
Reflexion (Shinn et al., 2023) QA self-critique $\checkmark$
RCI (Kim et al., 2023) Code, QA self-critique $\checkmark$
Tree Search Uninformed search ToT (Yao et al., 2023) Planing, Creative writing self-critique $\checkmark$
@@
Heuristic search RAP (Hao et al., 2023) Planing, Math, Logical self-critique $\checkmark$
TS-LLM (Feng et al., 2024a) Planing, Math, Logical ORM
rStar (Qi et al., 2024a) Math, QA multi-agent consistency $\checkmark$
ReST-MCTS* (Zhang et al., 2024m) Math, QA PRM
Table 2. Overview of search strategies.
4.2. Search Strategies
4.2.1. Repeated Sampling